Hopefully I put this in the correct category:
A right triangle is a triangle with one angle measuring 90°. In a right triangle, the lengths of the sides are related by the Pythagorean Theorem, , where c is the length of the hypotenuse (the side opposite the 90° angle). Find the length of the hypotenuse when the other two sides’ measurements are 6 feet and 8 feet.

Answer by checkley75(3666)

6^2+8^2=C^2
36+64=C^2
C^2=100
C=SQRT100
C=10 FOR THE HYPOTENUSE.